# Long-range Entanglement and Role of Realistic Interaction in Braiding of Non-Abelian Quasiholes in Fractional Quantum Hall Phases

### v2

@@ Abstract
Parity conservation dictates that when fusing pairs of Moore-Read (MR) quasiholes, such that each pair of charge-$e/4$ anyons forms a charge-$e/2$ anyon, the parity of the numbers of $1$-anyon and $\psi$-anyon must be conserved within a given system. This idea is illustrated here using the Jack polynomial formalism, which also provides a basis to numerically study the dynamics of MR anyons. We find that parity conservation gives rise to a long-range “entanglement” between statistics of one pair of anyons and the fusion channels of other anyons in the background. It is therefore important to account for all quasiholes in an experimental system in order to accurately predict the measurement outcome. We also examine the effect of two-body electron-electron interaction on the degeneracy of two anyon fusion channels. This understanding of the quasihole dynamics can help to fine-tune two-body interactions in order to stabilize the non-Abelian degeneracy, which is important for detecting non-Abelian braiding in experiments.
@@ I Introduction
All different models of anyons share a similarity: one anyon is not a singular particle, but rather a low-energy collective excitation. Ideally, they are realized by model Hamiltonians that are not readily accessible in real experiments. For example, Ising anyons in FQH phase are realized by the Moore-Read (MR) quasiholes at filling factor $\nu=5/2$[39, 40, 41, 42, 43, 12], which are exact zero-energy ground states of a three-body electron-electron interaction[42, 44]. While it is believed that the $5/2$ plateau observed in the experiment is a non-Abelian phase, experiments probing the statistics of its quasiholes are complicated by the quasihole dynamics arising from Coulomb-type two-body interactions[38]. Previous studies[44, 45, 46, 47, 35, 48] have highlighted several effects of Coulomb interaction on MR anyons: the splitting of the degeneracy required for non-Abelian statistics, the stability of the two fusion channels, etc. These studies, however, exhibit many caveats. For one, evaluating the Coulomb variational energy using the model wavefunction gives the physics in the lowest Landau level (LLL), whereas the MR state in experiments is realized on the second Landau level (1LL), whose properties have been shown to differ drastically from those of the LLL[25, 49, 50]. Furthermore, many past studies, often constrained by numerical calculations, are limited in their scopes – they study the energy spectra for only a handful of real-space configurations of anyons[44, 48], which may not a reflection of the full anyon dynamics that vary as a non-monotonous function of their pair-wise separation[50, 51]. An overall description of the microscopic mechanism for multiple fusion channels in non-Abelian anyons and the effects of realistic interaction is in general lacking.
@@
Anyon fusion is the central rule in determining all physical properties of an anyon model, including mutual statistics[26, 48]. In the context of FQH, each anyon is a quasihole excitation [^1] A quasielectron in a FQH system is also anyonic, but our study does not concern quasielectrons because they require a finite energy even with respect to model Hamiltonian., parametrized by a location and possibly a quantum metric describing its shape[53, 54, 55, 56, 57]. “Fusing” two anyons simply means bringing their positions together so that they can be “stacked” one on top of the other. In the context of Abelian braiding, this description gives rise to a microscopic derivation of the spin-statistics relation for anyons[57, 56]. In non-Abelian braiding, such as the MR phase, fusing two anyons may result in more than one outcome[34]. The (approximate) energetic degeneracy of these two fusion channels under realistic electron-electron interaction is then necessary for non-Abelian braiding[45, 44, 58, 38, 37]. In the MR state the dynamics of two- and four-quasihole systems have been numerically studied in two limits: when quasiholes are very far from each other[44, 59], and when two quasiholes completely fuse together[45]. However, at finite separation the energy gap may be finite as this degeneracy is not protected by any topology or symmetry. It is thus an important question as to how the electron-electron interaction can be tuned to close this gap and enable a platform for non-Abelian braiding, in systems where many quasiholes can be present and interacting with each other.
@@
Apart from energetics, the fusion channels of MR quasiholes are potentially dictated by another factor called parity conservation[48, 60]. As will be explained in detail in the next section, two MR quasiholes can fuse in two different channels, labeled “$1$” and “$\psi$”. The parity conservation places a constraint on the fusion outcome of the entire system, saying that the number of $\psi$ anyons in a given system always remains either even or odd. Aspects of this parity that have been studied include different edge mode countings in the “odd-sector” and “even-sector”[61], as well as effects of fusion channels on Abelian braiding in the case of two quasiholes[48, 38]. However, an important implication of the parity conservation that has not been pointed out and studied extensively is how the braiding of a set of anyons is influenced by not only their mutual statistics, but also the presence and fusion of any other anyons in the background even if the background anyons are very far away (see Fig.1).
@@
This paper is organized as follows. In the next section we review the Ising anyon model and illustrate, using the Jack polynomial formalism[65, 66], how the three types of Ising anyons manifest physically as MR quasiholes. We will also show that parity conservation can be seen from the root configurations of the Jack polynomials. Next, in Section III we discuss an important implication of parity conservation, which is a “long-range entanglement” that enables the fusion of one pair of anyons to directly affect the statistics of another pair elsewhere in the same system. In section IV, we numerically investigate the effects of two-body interactions on the energetics of MR quasiholes. This allows us to simulate the entanglement effect by demonstrating how simultaneous selection of fusion channel of two quasiholes in a four-quasihole MR system renders it Abelian with pre-determined statistics. We will also discuss methods to mitigate this effect to stabilize non-Abelian signatures in experiments.
@@ II.1 Types of MR quasiholes
In the MR state, $\sigma$-anyons are the charge-$e/4$ quasiholes, and $1$- and $\psi$-anyons are different types of charge-$e/2$ quasiholes of different types. Here we explicitly construct the microscopic wavefunctions for these three types of anyons using the Jack polynomial formalism. [65, 66]. This approach exhibits two major advantages, one conceptual and one practical. Conceptually, it provides a simple understanding of the parity conservation in terms of “root configurations”. Practically, the Jack polynomials provide a basis for obtaining via ED the low-lying states of Hamiltonian of the form given by Eq. (1), in the case where $\hat{H}_{\text{model}}=\hat{V}_{3}^{3bdy}$ is the three-body interaction and $\lambda_{0}\to\infty$. This enables us to greatly reduce the dimension of the Hilbert space involved, but with a small price that the Hamiltonian matrix is no longer sparse[44]. Throughout this paper, we work on the spherical geometry[67].
Figure 2: (a) The spherical geometry used in this paper[67]: a magnetic monopole of strength $2S\hbar$ (where $S$ is either integer or half-integer) is placed at the center of a sphere, creating a uniform magnetic field everywhere perpendicular to the surface of the sphere. The Landau orbitals are concentric rings parallel to the equator[68](b)The density per orbital on the sphere showing the three types of anyons in the Moore-Read state: two $\sigma$-anyons at each pole (blue); one $1$-type at the north pole (orange); and one $\psi$-type at the north pole (green).
Following the Jack polynomial formalism, the ground state and quasihole wavefunctions of the Moore-Read state are the Jack polynomials with root configurations that satisfies the “$(2,4)$-admissibility rule”. A root configuration can be represented by a string of digits 0’s and 1’s, and the $(2,4)$-admissible rule says that any group of four consecutive digits contains no more than two digit 1’s. Physically, each digit is interpreted as a Landau orbital, where “1” means an electron is occupying the corresponding orbital and “0” means the orbital is empty. On the spherical geometry, the left end of the root configuration also corresponds to the orbital at the north pole while the right end corresponds to the south pole (see Fig. 2). As a result of the $(2,4)$-admissible rule, a root with $N_{e}$ electrons must have at least $22N_{e}-$ 2 orbitals in total. There exists a unique root with $N_{e}$ electrons and $N_{o}=2N_{e}-2$ orbitals:
$1100110011...110011$ (5)
where “…” denotes a pattern of repeating “$1100$”. This root correspond to the MR ground state, whose first-quantized wavefunction is the well-known Pfaffian:
@@
Quasiholes can be added to the ground state by insertion of additional orbitals. Symbolically this is done by adding 0’s to the root configuration. One example is:
[…]
where we use the empty circles below the root configuration to denote the positions of the quasiholes. Quasiholes are located in the root configuration where four consecutive orbitals contain fewer than two electrons[69]. The above root configuration can be rearranged such that the two quasiholes are separated:
[…]
where the ellipsis now denotes repeating patterns of “10”. There exists another configuration where two quasiholes are closed together, such that they form a bound pair:
[…]
@@
[…]
If one applies a single one-body potential pin at the north pole on a MR system with an odd number of electrons (denoted as $N_{e}$) and $12N_{e}-$ 1 orbitals, then the ground state will be exactly the Jack given by root configuration in Eq.(10) and not Eq.(9). Since our analysis in this paper relies on the interplay between one-body and two-body interaction, we will use Eq.(10) instead of Eq.(9) as the model state for the $\psi$-anyon. The electron densities of the wavefunctions on the sphere corresponding to the three types of anyons are shown in Fig.2.
@@ II.2 Parity conservation in systems with four or more quasiholes
[…]
where in Eq.(13), we changed the empty circles into crosses to denotes the pair of quasiholes that fuses into a $\psi$-type anyon [^2] Two MR quasiholes are denoted as empty circles if there are an even number of digit 1’s between them, and as crosses if there are an odd number of 1’s between them.. This change of notation serves to distinguish different anyon species within the same system and makes it easier to observe the parity conservation. In this system with two pair of quasiholes, one can read off the fusion result of each pair of quasiholes (two consecutive empty circles) based on the number of electrons (digit 1’s) between them. Two quasiholes fuse into a $1$-anyon if there are an even number of electrons between them, and $\psi$-anyon if there are an odd number of electrons between them – this can also be observed in Eqs. (7) and (10). In Eqs. (11) and (12) there are two $1$-anyons and in Eqs. (13) and (14) there are two $\psi$-anyons. One can now observe that in a given root configuration a fixed number of orbitals and electrons, two $1$-anyons may be transmuted into two $\psi$-anyons and vice versa by shifting the electrons around, but the parity of the number of $\psi$-anyons (its oddness or evenness) is a conserved quantity.
@@
[…]

### v3

@@ Abstract
Parity conservation dictates that when fusing pairs of Moore-Read (MR) quasiholes, such that each pair of charge-$e/4$ anyons forms a charge-$e/2$ anyon, the parity of the numbers of $1$-anyon and $\psi$-anyon must be conserved within a given system. This idea is illustrated here using the Jack polynomial formalism, which also provides a basis to numerically study the dynamics of MR anyons. We find that parity conservation gives rise to a long-range ``entanglement" between statistics of one pair of anyons and the fusion channels of other anyons in the background. It is therefore important to account for all quasiholes in an experimental system in order to accurately predict the measurement outcome. We also examine the effect of two-body electron-electron interaction on the degeneracy of two anyon fusion channels. This understanding of the quasihole dynamics can help to fine-tune two-body interactions in order to stabilize the non-Abelian degeneracy, which is important for detecting non-Abelian braiding in experiments.
@@ I Introduction
All different models of anyons share a similarity: one anyon is not a singular particle, but rather a low-energy collective excitation. Ideally, they are realized by model Hamiltonians that are not readily accessible in real experiments. For example, Ising anyons in FQH phase are realized by the Moore-Read (MR) quasiholes at filling factor $\nu=5/2$[39, 40, 41, 42, 43, 12], which are exact zero-energy ground states of a three-body electron-electron interaction[42, 44]. While it is believed that the $5/2$ plateau observed in the experiment is a non-Abelian phase, experiments probing the statistics of its quasiholes are complicated by the quasihole dynamics arising from Coulomb-type two-body interactions[38]. Previous studies[44, 45, 46, 47, 35, 48] have highlighted several effects of Coulomb interaction on MR anyons: the splitting of the degeneracy required for non-Abelian statistics, the stability of the two fusion channels, etc. These studies, however, exhibit many caveats. For one, evaluating the Coulomb variational energy using the model wavefunction gives the physics in the lowest Landau level (LLL), whereas the MR state in experiments is realized on the second Landau level (1LL), whose properties have been shown to differ drastically from those of the LLL[25, 49, 50]. Furthermore, many past studies, often constrained by numerical calculations, are limited in their scopes – they study the energy spectra for only a handful of real-space configurations of anyons[44, 48], which may not be a reflection of the full anyon dynamics that vary as a non-monotonous function of their pair-wise separation[50, 51]. An overall description of the microscopic mechanism for multiple fusion channels in non-Abelian anyons and the effects of realistic interaction is in general lacking.
@@
Anyon fusion is the central rule in determining all physical properties of an anyon model, including mutual statistics[26, 48]. In the context of FQH, each anyon is a quasihole excitation [^1] A quasielectron in a FQH system is also anyonic, but our study does not concern quasielectrons because they require a finite energy even with respect to model Hamiltonian., parametrized by a location and possibly a quantum metric describing its shape[53, 54, 55, 56, 57]. ``Fusing" two anyons simply means bringing their positions together so that they can be ``stacked" one on top of the other. In the context of Abelian braiding, this description gives rise to a microscopic derivation of the spin-statistics relation for anyons[57, 56]. In non-Abelian braiding, such as the MR phase, fusing two anyons may result in more than one outcome[34]. The (approximate) energetic degeneracy of these two fusion channels under realistic electron-electron interaction is then necessary for non-Abelian braiding[45, 44, 58, 38, 37]. In the MR state the dynamics of two- and four-quasihole systems have been numerically studied in two limits: when quasiholes are very far from each other[44, 59], and when two quasiholes completely fuse together[45]. However, at finite separation the energy gap may be finite as this degeneracy is not protected by any topology or symmetry. It is thus an important question as to how the electron-electron interaction can be tuned to close this gap and enable a platform for non-Abelian braiding, in systems where many quasiholes can be present and interacting with each other.
@@
Apart from energetics, the fusion channels of MR quasiholes are potentially dictated by another factor called parity conservation[48, 60]. As will be explained in detail in the next section, two MR quasiholes can fuse in two different channels, labeled ``$1$'' and $\psi''$. The parity conservation places a constraint on the fusion outcome of the entire system, saying that the number of $\psi$ anyons in a given system always remains either even or odd. Aspects of this parity that have been studied include different edge mode countings in the ``odd-sector" and ``even-sector"[61], as well as effects of fusion channels on Abelian braiding in the case of two quasiholes[48, 38]. However, an important implication of the parity conservation that has not been pointed out and studied extensively is how the braiding of a set of anyons is influenced by not only their mutual statistics, but also the presence and fusion of any other anyons in the background even if the background anyons are very far away (see Fig.1).
@@
This paper is organized as follows. In the next section we review the Ising anyon model and illustrate, using the Jack polynomial formalism[65, 66], how the three types of Ising anyons manifest physically as MR quasiholes. We will also show that parity conservation can be seen from the root configurations of the Jack polynomials. Next, in Section III we discuss an important implication of parity conservation, which is a ``long-range entanglement'' that enables the fusion of one pair of anyons to directly affect the statistics of another pair elsewhere in the same system. In section IV, we numerically investigate the effects of two-body interactions on the energetics of MR quasiholes. This allows us to simulate the entanglement effect by demonstrating how simultaneous selection of fusion channel of two quasiholes in a four-quasihole MR system renders it Abelian with pre-determined statistics. We will also discuss methods to mitigate this effect to stabilize non-Abelian signatures in experiments.
@@ II.1 Types of MR quasiholes
In the MR state, $\sigma$-anyons are the charge-$e/4$ quasiholes, and $1$- and $\psi$-anyons are charge-$e/2$ quasiholes of different types. Here we explicitly construct the microscopic wavefunctions for these three types of anyons using the Jack polynomial formalism. [65, 66]. This approach exhibits two major advantages, one conceptual and one practical. Conceptually, it provides a simple understanding of the parity conservation in terms of ``root configurations". Practically, the Jack polynomials provide a basis for obtaining via ED the low-lying states of Hamiltonian of the form given by Eq. (1), in the case where $\hat{H}_{\text{model}}=\hat{V}_{3}^{3bdy}$ is the three-body interaction and $\lambda_{0}\to\infty$. This enables us to greatly reduce the dimension of the Hilbert space involved, but with a small price that the Hamiltonian matrix is no longer sparse[44]. Throughout this paper, we work on the spherical geometry[67].
Figure 2: (a) The spherical geometry used in this paper[67]: a magnetic monopole of strength $2S\hbar$ (where $S$ is either integer or half-integer) is placed at the center of a sphere, creating a uniform magnetic field everywhere perpendicular to the surface of the sphere. The Landau orbitals are concentric rings parallel to the equator[68](b) A root configuration is a binary string with each digit representing an orbital from the north pole to the south pole going right to left. To read off the position of MR quasiholes, every group of four consecutive digits is checked (examplified by yellow and purple ellipses) and a quasihole is marked whenever there are fewer than two electrons in it (examplified by yellow ellipses). The root is extended on the two sides by ``virtual'' repetitions of ``1100'' in order to read off the quasiholes residing at the two poles. (c) The density per orbital on the sphere showing the three types of anyons in the Moore-Read state: two $\sigma$-anyons at each pole (blue); one $1$-type at the north pole (orange); and one $\psi$-type at the north pole (green).
Following the Jack polynomial formalism, the ground state and quasihole wavefunctions of the Moore-Read state are the Jack polynomials with root configurations that satisfies the ``$(2,4)$-admissibility rule". A root configuration can be represented by a string of digits 0's and 1's, and the $(2,4)$-admissible rule says that any group of four consecutive digits contains no more than two digit 1's. Physically, each digit is interpreted as a Landau orbital, where ``1" means an electron is occupying the corresponding orbital and ``0" means the orbital is empty. On the spherical geometry, the left end of the root configuration also corresponds to the orbital at the north pole while the right end corresponds to the south pole (see Fig. 2). As a result of the $(2,4)$-admissible rule, a root with $N_{e}$ electrons must have at least $2N_{e}-2$ orbitals in total. There exists a unique root with $N_{e}$ electrons and $N_{o}=2N_{e}-2$ orbitals:
11001100111100111100110011...110011 (5)
where ``…" denotes a pattern of repeating ``11001100". This root correspond to the MR ground state, whose first-quantized wavefunction is the well-known Pfaffian:
@@
Quasiholes can be added to the ground state by insertion of additional orbitals. Symbolically this is done by adding 0's to the root configuration. One example is:
 1 $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}01100110011...110011$ (7)
where we use the empty circles below the root configuration to denote the positions of the quasiholes. Quasiholes are located in the root configuration where four consecutive orbitals contain fewer than two electrons (see Fig. 2b)[69]. Note that while this is seemingly a tongue-in-cheek method to mark down quasihole positions, there exists an empirical correspondence between the orbital positions marked in the root configuration and the deficiency in electron density in the actual state obtained from the Jack polynomial parametrized by that root[70, 71, 50]. This is a special aspect that makes the Jack formalism especially useful in studying the physics of quasiholes[50, 51].
The 0's and 1's in the root configuration shown in Eq. 7 can be shuffled such that the two quasiholes are separated:
 11 $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-5.7083pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ{\color[rgb]{1,1,1}{1}}$}}\hskip-5.7083pt\hskip 2.5pt}010101...1010\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-6.21524pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu{\color[rgb]{1,1,1}{1}}\scriptstyle\circ$}}\hskip-6.21524pt\hskip 2.5pt}$ (8)
where the ellipsis now denotes repeating patterns of ``10". There exists another configuration where two quasiholes are closed together, such that they form a bound pair:
 1 $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}00110011...11001$1 (9)
@@
 11 010 11 $\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-5.7083pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ{\color[rgb]{1,1,1}{1}}$}}\hskip-5.7083pt\hskip 2.5pt}010\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-6.21524pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu{\color[rgb]{1,1,1}{1}}\scriptstyle\circ$}}\hskip-6.21524pt\hskip 2.5pt}00110011...1$10011 (10)
If one applies a single one-body potential pin at the north pole on a MR system with an odd number of electrons (denoted as $N_{e}$) and $2N_{e}-1$ orbitals, then the ground state will be exactly the Jack given by root configuration in Eq.(10) and not Eq.(9). Since our analysis in this paper relies on the interplay between one-body and two-body interaction, we will use Eq.(10) instead of Eq.(9) as the model state for the $\psi$-anyon. The electron densities of the wavefunctions on the sphere corresponding to the three types of anyons are shown in Fig.2.
@@ II.2 Parity conservation in systems with four or more quasiholes
 1  1  0 $1\displaystyle\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-3.71524pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ$}}\hskip-3.71524pt\hskip 2.5pt}\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-3.71524pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle\circ$}}\hskip-3.71524pt\hskip 2.5pt}\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}011001100110011...110$011 (11)
 1 0110 0 $11\displaystyle\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}0110\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$0$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}01100110$011...110011 (12)
 1×1 010 11× $0$ 1×× $1\displaystyle\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-7.18748pt\raisebox{-4.7222pt}{\hbox{$\mkern 2.0mu\scriptscriptstyle\times\color[rgb]{1,1,1}{1}$}}\hskip-7.18748pt\hskip 2.5pt}010\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-6.875pt\raisebox{-4.7222pt}{\hbox{$\mkern 2.0mu{\color[rgb]{1,1,1}{1}}\scriptscriptstyle\times$}}\hskip-6.875pt\hskip 2.5pt}00\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-5.97227pt\raisebox{-4.7222pt}{\hbox{$\mkern 2.0mu\scriptscriptstyle{\times\times}$}}\hskip-5.97227pt\hskip 2.5pt}0011$0011...110011 (13)
@@
$\vdots\hskip 56.9055pt$
where in Eq.(13), we changed the empty circles into crosses to denotes the pair of quasiholes that fuses into a $\psi$-type anyon [^2] Two MR quasiholes are denoted as empty circles if there are an even number of digit 1’s between them, and as crosses if there are an odd number of 1’s between them.. This change of notation serves to distinguish different anyon species within the same system and makes it easier to observe the parity conservation. In this system with two pair of quasiholes, one can read off the fusion result of each pair of quasiholes (two consecutive empty circles) based on the number of electrons (digit 1's) between them. Two quasiholes fuse into a $1$-anyon if there are an even number of electrons between them, and $\psi$-anyon if there are an odd number of electrons between them – this can also be observed in Eqs. (7) and (10). In Eqs. (11) and (12) there are two $1$-anyons and in Eqs. (13) and (14) there are two $\psi$-anyons. One can now observe that in a given root configuration a fixed number of orbitals and electrons, two $1$-anyons may be transmuted into two $\psi$-anyons and vice versa by shifting the electrons around, but the parity of the number of $\psi$-anyons (its oddness or evenness) is a conserved quantity.
@@
 1 0 $1\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$\color[rgb]{1,1,1}{1}$}\hskip-2.5pt\hskip-4.65277pt\raisebox{-5.55556pt}{\hbox{$\mkern 2.0mu\scriptstyle{\circ\circ}$}}\hskip-4.65277pt\hskip 2.5pt}0\hbox{\hskip 2.5pt\hskip-2.5pt\hbox{$1$}\hskip-2.5pt\hskip-5.97227pt\raisebox{-4.7222pt}{\hbox{$\mkern 2.0mu\scriptscriptstyle{\times\times}$}}\hskip-5.97227pt\hskip 2.5pt}00110011...110$011 (17)